# Cell as Point: One-Stage Framework for Efficient Cell Tracking

### v2

@@ Abstract
Conventional multi-stage cell tracking approaches rely heavily on detection or segmentation in each frame as a prerequisite, requiring substantial resources for high-quality segmentation masks and increasing the overall prediction time. To address these limitations, we propose CAP, a novel end-to-end one-stage framework that reimagines cell tracking by treating Cell as Point. Unlike traditional methods, CAP eliminates the need for explicit detection or segmentation, instead jointly tracking cells for sequences in one stage by leveraging the inherent correlations among their trajectories. This simplification reduces both labeling requirements and pipeline complexity. However, directly processing the entire sequence in one stage poses challenges related to data imbalance in capturing cell division events and long sequence inference. To solve these challenges, CAP introduces two key innovations: (1) adaptive event-guided (AEG) sampling, which prioritizes cell division events to mitigate the occurrence imbalance of cell events, and (2) the rolling-as-window (RAW) inference strategy, which ensures continuous and stable tracking of newly emerging cells over extended sequences. By removing the dependency on segmentation-based preprocessing while addressing the challenges of imbalanced occurrence of cell events and long-sequence tracking, CAP demonstrates promising cell tracking performance and is 10 to 55 times more efficient than existing methods. The code and model checkpoints will be available soon.
@@ 1 Introduction
Figure 1: Our proposed CAP is an end-to-end trained framework leveraging the idea of Cell as Point to facilitate tracking cells efficiently. As (a) illustrates, instead of previous work requiring segmentation (SEG) or detection (DET) as a prerequisite for final tracking (TRA), CAP is able to track all cells within sequence frames in one stage. (b) shows that CAP reduces the inference time ($2.9$s) by approximately $\mathbf{10}$ to $\mathbf{55}$ times compared to previous works while maintaining high tracking performance of $0.93$. (c) demonstrates the cell tracking result predicted by CAP.
@@ 2.1 Cell Tracking
Cell tracking is a crucial task in biological imaging that typically involves the identification and temporal association of cell instances across image frames. Traditionally, this process includes two main steps: cell segmentation (detection) [1, 79, 55, 77] and cell linking [18]. The segmentation step identifies individual cells in each frame, while the linking step establishes temporal associations between these cells across frames, sometimes capturing complex dynamics such as cell division and migration [49, 52]. Nowadays, instead of traditional approaches such as Nearest Neighbors linking [67], Bayesian filters [66, 16, 33, 81, 34, 73], Viterbi [51, 50], and graph-based matching [5, 72, 3, 48], the advent of deep learning has transformed the landscape of cell tracking, leveraging advanced neural architectures for segmentation and tracking tasks. Deep learning approaches often employ complex models, e.g., CNN, RNN, GNN, and transformer, to combine multiple stages like detection [12, 10, 11, 46, 88, 30, 56, 78, 31, 32, 38], segmentation [82, 59, 9, 2, 49, 68, 7, 8, 64, 58, 17, 13, 57, 14, 71], and tracking into the frameworks. Several contemporary methods [62, 60, 30] aim to integrate contextual cell data into tracking models. ConvGRU [60] is introduced by Payer et al., capturing local cellular and inter-frame features. However, this method requires extensive annotated training data for all cell regions, making it labor-intensive. Motion and Position Map (MPM) is capable of representing detection and association at once, which is proposed by Hayashida et al. [31], marking an advancement in simplifying the pipeline for cell tracking. However, it still requires multiple stages, first generating the MPM followed by subsequent detection and tracking procedures.
@@ 2.2 Point Tracking
3 Cell as Point (CAP)
Figure 2: Cell Point Trajectory and Visibility. M, D1, and D2 denote mother cell, daughter cell #1, and daughter cell #2. (x,y) in $L_{t}$ represent the location coordinates, and 00 or $1$ in $V_{t}$ respectively represent non-existing and existing cells. Three valid trajectory and visibility combinations: (a) M has not divided; (b) M has divided, and only D1 occurs in the $t$-th frame; (c) M has divided, and both D1 and D2 occur in the $t$-th frame.
@@ 3 Cell as Point (CAP)
3.1 Adaptive Event-Guided (AEG) Sampling
@@
3.2 Cell Joint Tracking
@@
To initialize the tracking feature $F$ for the part of input tokens $G$, we first generate image features (feature map), $\phi(I_{t})$, for each frame in the training sequence, from which the tracking features, $F$, are distilled according to cell point locations. Specifically, the feature extractor is a convolutional neural network (CNN) trained end-to-end in the framework for computational efficiency. Leveraging the CNN feature extractor, the feature maps with $dim$ dimensions are generated from $I_{t}$, $\phi(I_{t})\in\mathbb{R}^{dim\times\left(\left\lfloor\frac{H}{k}\right\rfloor\times\left\lfloor\frac{W}{k}\right\rfloor\right)}$, where $k=4$ for computational efficiency, and the features are downscaled by average pooling with strides $s$ for multi-scaled features. The tracking features $F^{i}_{t=1}$ for $i$ are captured by bilinear interpolation from the location of cell points on the initial frame $L_{t=1}$ on the feature map $\phi(I_{t=1})$. The tracking feature vector $F^{i}_{t}\in\mathbb{R}^{dim}$ is initialized for the transformer input by broadcasting the tracking features in $t=1$ frame for all $T_{s}$ frames, which is refined and updated with $M$ iterations.
@@
where $s=1,\dots,S$ are the feature scales, and $\delta$ is the offset around cell points, $(\hat{x},\hat{y})$ is the estimated location coordinate of the cell point $\hat{L}_{t}^{i}$. The value of $[C_{t}^{i}]_{\delta}^{s}$ is computed as the inner product between the feature vector $F^{i}_{t}$ and $\phi_{s}(I_{t})$ at $\hat{L}_{t}^{i}$ pooled with kernel size $2^{s}\times 2^{s}$.
@@
3.3 Rolling-as-Window (RAW) Inference
Metrics PhC-C2DH-U373 Fluo-N2DH-GOWT1 Fluo-C2DL-Huh7
GNN EmbedTrack KIT-GE CAP GNN EmbedTrack KIT-GE CAP GNN EmbedTrack KIT-GE CAP
$\mathrm{ES}$downarrow 214 0 3 0 49 39 45 18 34 0 34 31
$\mathrm{EA}$downarrow 371 44 50 0 140 119 111 120 85 733 77 71
$\mathrm{ED}$downarrow 0 0 0 0 1 0 0 2 2 0 0 0
$\mathrm{ESM}$downarrow 0 7 1 0 13 0 1 2 835 5 0 1
$\mathrm{FP}$downarrow 135 211 197 119 5 21 6 75 36 1392 41 101
$\mathrm{FN}$downarrow 25 32 25 0 28 40 32 46 4 636 0 1
$\mathrm{TRA}$uparrow 0.771 0.931 0.939 0.985 0.958 0.961 0.966 0.960 0.889 0.187 0.970 0.966
$\mathrm{IT}$downarrow 19.7 123.6 22.0 2.6 67.2 171.6 39.4 7.1 27.1 125.0 25.7 3.2
Metrics HeLa PC-3 3T3 RAW264
GNN EmbedTrack CAP GNN EmbedTrack CAP GNN EmbedTrack CAP GNN EmbedTrack CAP
$\mathrm{ES}$downarrow 2 0 3 2 0 1 6 2 107 186 68 67
$\mathrm{EA}$downarrow 385 211 82 52 43 11 613 603 949 1451 1475 1098
$\mathrm{ED}$downarrow 1 1 1 3 3 0 16 14 2 89 74 12
$\mathrm{ESM}$downarrow 13 1 0 12 1 0 3 4 4 21 28 17
$\mathrm{FP}$downarrow 62 31 60 33 0 35 100 201 250 239 669 647
$\mathrm{FN}$downarrow 342 218 70 0 42 7 9 589 709 1341 1259 898
$\mathrm{TRA}$uparrow 0.776 0.849 0.926 0.868 0.912 0.952 0.857 0.853 0.832 0.801 0.803 0.864
$\mathrm{IT}$downarrow 23.3 41.5 1.3 18.3 48.7 1.1 21.7 74.6 2.2 28.5 59.6 2.8
Table 1: Comparison Experiments with Baselines. $\mathrm{ES}$, $\mathrm{EA}$, $\mathrm{ED}$, and $\mathrm{ESM}$ are abbreviations of $\mathrm{EdgeSplit}$, $\mathrm{EdgeAdd}$, $\mathrm{EdgeDelete}$, and $\mathrm{EdgeSemantic}$, respectively. Our framework CAP is TRA only, while previous frameworks GNN, EmbedTrack, and KIT-GE are SEG+TRA illustrated in Fig. 1 (a). Due to space limitations, KIT-GE results on DeepCell datasets are included in the supplementary. Average inference time/seconds ($\mathrm{IT}$) and performance ($\mathrm{TRA}$) are visualized in Fig. 1 (b).
In cell tracking, the inference stage is crucial for producing the final tracking mask and establishing cell lineage relationships by analyzing predicted cell locations and division events. Existing works [68, 64, 78, 49, 8] often customize inference and post-processing stages with different configurations for each dataset, adapting to variations in sequence length. This data-specific tailoring limits their generalizability and broader applicability to other datasets. In contrast, our proposed Rolling-As-Window (RAW) inference approach in CAP simplifies and unifies these processes while maintaining effectiveness across varying sequence lengths, enabling high-quality inference regardless of computational constraints. Specifically, a consistent and high-quality inference result is achieved through a rolling loop, which sequentially extracts yet-to-be-inferred clips from the full sequence, processing them within a defined window for inference.
@@ 4.1 Experimental Setup
Figure 5: Cell Divisions. (a) and (b) illustrate cell division events accurately detected by the CAP framework in the HeLa and 3T3 datasets, respectively.
Datasets.
@@
We evaluate our proposed framework on DynamicNuclearNet Tracking in DeepCell [54] public database, HeLa, PC-3, 3T3, and RAW264 datasets divided by specimen, and ISBI cell tracking challenge [52, 74] public datasets, PhC-C2DH-U373 (U373), Fluo-N2DH-GOWT1 (GOWT1), and Fluo-C2DL-Huh7 (Huh7). These temporal data used for cell tracking, shot by different devices in various durations on diverse specimens, exhibit diverse morphologies and sequence characteristics. DeepCell contains $96$ sequences in total: $45$ for HeLa, $7$ for PC-3, $10$ for 3T3, and $34$ for RAW264, with $42$, $50$, $71$, and $45$ frames in each sequence, respectively. For the ISBI dataset, the U373 comprises glioblastoma-astrocytoma U373 cells in a $115$-frame sequence of size $520\times 696$. Dataset GOWT1 documents GFP-GOWT1 mouse stem cells into $92$ frames with a size of $1024\times 1024$ per sequence. Huh7 is a dataset of human hepatocarcinoma-derived cells in a sequence of $30$ frames with a $1024\times 1024$ size.
Implementation Details.
For the training process, we utilize $85$ training sequences from the DeepCell dataset and sequences $02$ from the ISBI dataset, using only tracking (TRA) ground truth (GT). During each iteration, we crop sequences of $T_{s}=24$ frames with a sliding window length of $l_{\text{slide}}=8$. We use pre-trained checkpoints from CoTracker [40], which was trained solely on synthetic RGB data. Since our input images are grayscale, we replicate the single channel three times to match the required RGB format. We use an end-to-end trained $2$-dimensional CNN architecture as the encoder to extract tracking features $F$. Specifically, it integrates bilinear interpolation and utilizes it to extract $S=4$ different scaled image features with $128$-channel output features concatenated to optimized dynamic feature extraction (Sec. 3.2). While training the transformer, we follow the pre-trained checkpoint that downsamples the resolution of frames to $384\times 512$ and choose $S=4$, $\Delta=3$ to calculate 4D multi-scaled correlation volumes (Sec. 3.2). More details can be found in the supplementary.
As for the testing (evaluation) process, we preprocess the first frame for each sequence using [64] to select query points $Q$ and start the process of window rolling. The evaluation sequence $S_{\text{infer}}$ follows the native resolution ($384\times 512$), which the dataset is trained on. We set $l_{\text{win}}=100$, which is mentioned in Sec. 3.3. We use the $\mathrm{TRA}$ evaluation criteria provided by ISBI [52, 74] and CTMC [4], which lay in the range $[0,1]$ that a higher score corresponds with a better performance and is calculated by $\mathrm{TRA}=1-\frac{\mathrm{min}(AOGM,AOGM_{0})}{AOGM_{0}}$, where $AOGM$ is the measure of acyclic oriented graphs matching [53], penalizing the number of transformations needed to transform the predicted tracking graph into the ground-truth tracking graph, e.g., merged cells ($\mathrm{EdgeSplit}$), missing links ($\mathrm{EdgeAdd}$), wrong links ($\mathrm{EdgeDelete}$), links with wrong semantics ($\mathrm{EdgeSemantic}$), false positives ($\mathrm{FP}$), and false negatives ($\mathrm{FN}$), and $AOGM_{0}$ is the $AOGM$ value for empty tracking results.
Baselines.
We compare our approach CAP with three methods, cell-tracker-gnn (GNN) [8], EmbedTrack [49], and KIT-GE [64], which have publicly available code and utilize the evaluation system provided by ISBI [52, 74] and acyclic oriented graphs matching (AOGM) measure [53]. For the ISBI dataset, to control for variables and ensure fairness, we train these methods on sequence $02$ and evaluate on $01$, following [64]. For the DeepCell data, we follow the experiments [54] and report the result. We report the measures involved in the cell tracking, $\mathrm{EdgeSplit}$, $\mathrm{EdgeAdd}$, $\mathrm{EdgeDelete}$, $\mathrm{EdgeSemantic}$, $\mathrm{FP}$, $\mathrm{FN}$, $\mathrm{TRA}$, and average inference time ($\mathrm{IT}$).
4.2 Results and Analysis
Tab. 1 presents the quantitative performance comparison of our method, CAP, and baseline methods, GNN, EmbedTrack, and KIT-GE on the ISBI dataset (U373, GOWT1, and Huh7) and DeepCell dataset (HeLa, PC-3, 3T3, and RAW264). Since $\mathrm{TRA}$ is derived from metrics $\mathrm{EdgeSplit}$, $\mathrm{EdgeAdd}$, $\mathrm{EdgeDelete}$, $\mathrm{EdgeSemantic}$, $\mathrm{FP}$, and $\mathrm{FN}$, we primarily focus on the $\mathrm{TRA}$ score for the model performance and $\mathrm{IT}$ for efficiency (highlighted in gray).
U373 dataset with a $115$-frame sequence not only does CAP outperform all GNN, KIT-GE, and EmbedTrack significantly, achieving a perfect score of 00 for $\mathrm{EdgeSplit}$, $\mathrm{EdgeAdd}$, $\mathrm{EdgeDelete}$, $\mathrm{EdgeSemantic}$, and $\mathrm{FN}$ with the highest $\mathrm{TRA}$ score of $0.985$, but also it has the smallest inference time of $2.6$s. In contrast, GNN, EmbedTrack, and KIT-GE exhibit higher error counts and inference time. The $\mathrm{TRA}$ of datasets HeLa, PC-3, and RAW264, $0.926$, $0.952$, and $0.864$, show that The CAP not only significantly leads other methods in terms of the $\mathrm{TRA}$ values but also utilizes markedly less inference time compared to other approaches. CAP still demonstrates comparable performance with the $\mathrm{TRA}$ score of $0.960$, $0.966$, and $0.832$ across previous works in datasets GOWT1, Huh7, and 3T3. The trade-off remains advantageous given that CAP directly eliminates the stages of segmentation or detection, thereby significantly reducing the volume of data labels required and inference time ($7.1$s, $3.2$s, and $2.2$s) by approximately $5.5$ to $39$ times, while it has a slightly lower $\mathrm{TRA}$ measure than the existing method. Fig. 4 highlights CAP’s robust cell trajectory tracking in time-lapse microscopy, while Fig. 5 demonstrates its strong capability in detecting cell divisions, accurately tracking mother and daughter cells across frames.
As for the extreme instability of EmbedTrack method on Huh7 with $\mathrm{TRA}=0.187$, our analysis suggests that the observed results stem from the method’s excessive reliance on the segmentation (SEG) stage and the usage of refined masks during the training of the segmentation model, while Huh7 contains only SEG golden truth (GT) masks without the more refined SEG silver truth (ST). The supplementary Appendix B illustrates more details about the difference between tracking GT mask, SEG GT, and SEG ST. This further substantiates an end-to-end one-stage cell tracking framework, which bypasses the segmentation or detection stage, enhances model stability, and raises the overall performance limits achievable by the model.
Overall, the CAP method demonstrates robust performance and constant high efficiency on inference and label usage across all datasets, consistently achieving high $\mathrm{TRA}$ scores and the lowest $\mathrm{IT}$s. Thus, CAP is a highly effective, efficient, and reliable method for cell tracking in a practical biomedical context.
4.3 Ablation Study
     AEG Sampling      $\mathrm{TRA}$ Score
     U373      GOWT1      Huh7
     ✗      0.556      0.941      0.942
     ✓      0.985      0.960      0.966
Effectiveness of AEG Sampling.
To mitigate the data imbalance in cell division in samples, AEG sampling Sec. 3.1 is an effective method we proposed to calibrate the sample experience in a complete cell division process. Tab. 2 illustrates that switching off AEG sampling decreases performance from $0.02$ to $0.43$.
Attention $\mathrm{TRA}$ Score
time joint U373 GOWT1 Huh7
0.974 0.903 0.944
0.985 0.960 0.966
Effectiveness of Cell Joint Tracking.
By replacing the cross-trajectory attention layers with only time attention layers to keep the model size, Tab. 3 shows model performance declined by $0.015$ to $0.055$ across datasets. Thus, cell joint tracking plays an important role in CAP.
      Stride       $\mathrm{TRA}$ Score
      U373       GOWT1       Huh7
      8       0.531       0.609       0.853
      4       0.985       0.960       0.966
Ablation on Feature Stride.
Tab. 4 shows that the higher the resolution of the feature scaled, the more accurately CAP predicted. Therefore, we choose $s=4$ for scaling features during the training procedure (Sec. 3.2).
Additional ablations on DeepCell datasets are included in the supplementary.
5 Conclusion and Future Work
In this paper, we introduce the one-stage CAP (Cell as Point) framework, which eliminates the need for a separate segmentation or detection stage—a significant innovation in cell tracking. This approach substantially improves efficiency by reducing inference time by a factor of $10$ to $55$ and minimizing the need for high-quality data annotation. CAP effectively captures complex cell states and tracks cells in a single step, all while maintaining high efficiency. Additionally, it addresses common challenges in cell sequence data, such as cell event imbalance and long sequence inference. In conclusion, CAP not only offers considerable practical value in real-world applications due to its efficiency but also pioneers a novel approach in the field of cell tracking.
Limitations and Future Work.
Different datasets are captured by different devices, potentially resulting in variations in the luminance of cells presented in the time-lapse microscopy images; the datasets we utilized to evaluate our framework are annotated by various experts, so the quality of labels and annotation practices may vary; and different cell densities for datasets may cause disparities in performance. Consequently, a strategy to train a versatile model via our one-stage framework that could mitigate the gap and generalize to all datasets deserves future exploration.
Supplementary Material
This supplementary material complements the main manuscript by providing detailed information and additional support. It is structured as follows:
Appendix A: Additional Related Work
- Provides more related work about computing visual similarity and iterative updates proposed by RAFT.
Appendix B: Dataset Details
- Provides more detailed information and analysis on the ISBI and DeepCell datasets we utilized.
Appendix C: Experiment Details
- Offers the training and optimizer configuration details and evaluation metrics calculation details.
Appendix D: Additional Baseline Results
- Provides KIT-GE baseline method results on the DeepCell dataset as complementary for Tab. 1 in Sec. 4.2.
Appendix E: Additional Ablations
- Shows the effectiveness of adaptive event-guided sampling, cell joint tracking structure, and different feature strides in our proposed CAP framework.
Appendix A Additional Related Work
RAFT [70] is introduced by Teed and Deng, which further refined the DCFlow approach by incorporating incremental flow updates based on 4D correlation volumes constructed for computing the similarity between all pairs of pixels across two images [29], enabling the model to handle large displacements and complex motion patterns effectively. The model iteratively updates the flow field using a recurrent unit that performs lookups on the correlation volumes, leading to precise motion estimation [65].
A.1 Computing Visual Similarity
A.1.1 Correlation Volume
Given two feature maps $F_{1}\in\mathbb{R}^{H\times W\times dim}$ and $F_{2}\in\mathbb{R}^{H\times W\times dim}$ with output $dim$-dimensions, which are features extracted from two input images $I_{1}$ and $I_{2}$, by feature extractor $g_{\theta}$:
$F_{1}=g_{\theta}(I_{1})\text{ and }F_{2}=g_{\theta}(I_{2}).$ (6)
RAFT computes a 4D correlation volume $C$ that captures the similarity between all pairs of pixels across the two images. The correlation between a pixel at position $(i,j)$ in $F_{1}$ and a pixel at position $(k,l)$ in $F_{2}$ is calculated as the dot product of their feature vectors:
$C_{ijkl}$ $=\langle F_{1}(i,j),F_{2}(k,l)\rangle$ (7)
$=\sum_{h}F_{1}(i,j,h)\cdot F_{2}(k,l,h).$
The correlation volume is a 4D tensor of dimensions $H\times W\times H\times W$, where $H$ and $W$ are respectively the height and width of feature maps $F$s.
A.1.2 Correlation Pyramid
RAFT constructs a correlation pyramid by applying average pooling to the last two dimensions of the 4D correlation volume to capture multi-scale feature information. Specifically, this pooling is performed with kernel sizes of $1$ $2$, $4$, and $8$, and a set of multi-scale correlation volumes $C^{k}=\{C^{1},C^{2},C^{3},C^{4}\}$ is generated. Thus, each volume $C^{k}$ in the pyramid has dimensions $H\times W\times\frac{H}{2^{k}}\times\frac{W}{2^{k}}$.
A.2 Iterative Updates
The update operator estimates a sequence of flow estimations $\{f_{1},\dots,f_{n}\}$ with initializing the start point $f_{0}=0$. For each iteration, it produces $\Delta f$ as the direction that should be applied to the current estimation:
$f_{k+1}=\Delta f+f_{k+1}.$ (8)
Specifically, given the current flow estimation $f_{k}$, the correlation volume pyramid is retrieved and undergoes processing through two convolutional layers to extract relevant information. The updated information is comprised of correlation, flow, and context features and is updated over time as a feature map concatenation.
Appendix B Dataset Details
B.1 ISBI Dataset
Figure 7: Visualization of the DeepCell Dataset. HeLa, PC-3, 3T3, and RAW264 have different cell densities, sizes, and luminances.
To support the claim in Sec. 4.2, Fig. 6 illustrates datasets we leveraged, PhC-C2DH-U373, Fluo-N2DH-GOWT1, and Fluo-C2DL-Huh7 [52, 74], with the difference between tracking (TRA) GT mask, segmentation (SEG) GT, and segmentation (SEG) ST if available. No segmentation ST is available specific to dataset Fluo-C2DL-Huh7. We observe that SEG GT and ST are much more refined and high-quality than TRA GT masks, requiring more resources and manpower. Fig. 6 also demonstrates the comparison of data annotation requirements between the previous works and our framework CAP that, instead of previous works leveraging all TRA GT masks, segmentation GT, and ST if available, CAP utilizes the TRA GT masks. Thus, in addition to being a highly efficient framework for inference time, the CAP also excels in efficiency with respect to both the quality and quantity of labels required, which is attributed to the one-stage characteristic of the framework.
Furthermore, detailed information on data collection and the characteristics of each dataset are described below:
  • Fluo-C2DL-Huh7 is a dataset of human hepatocarcinoma-derived cells expressing the fusion protein YFP-TIA-1. The Nikon Eclipse Ti2 microscope with a lens CFI Plan Apo Lambda 20x/0.75 is the shooting device to capture the cell migration, death, and division in a duration of 15 minutes with $30$ frames for each sequence. Cells in this dataset exhibit a relatively low frequency of cell events; However, due to the smaller number of frames recorded within the same time period, some cells exhibit significantly larger movements than others within a sequence. Cells are relatively small and have high density in frames.
B.2 DeepCell Dataset
The DynamicNuclearNet Tracking in the DeepCell [54] database consists of HeLa, PC-3, 3T3, and RAW264 datasets, which respectively contains cell type HeLa-S3, PC-3, NIH-3T3, and RAW 264.7, visualized in Fig. 7. These four cell types are mammalian cell lines that were acquired from the American Type Culture Collection [54]. The cells are cultured in Dulbecco’s modified Eagle’s medium (RAW 264.7 and NIH-3T3) or F-12K medium (Hela-S3 and PC-3) supplemented with the necessary drugs, e.g., L-glutamine, penicillin, for all other cells.
There are $5$, $2$, $1$, and $3$ evaluation sequences for datasets HeLa, PC-3, 3T3, and RAW264, respectively. In the main manuscript, we report the average metric values for each dataset. In the following additional ablation study Appendix E in this supplementary, we report the detailed metric values for each sequence.
The detailed information on each cell type is illustrated below:

### v3

@@ Abstract
Conventional multi-stage cell tracking approaches rely heavily on detection or segmentation in each frame as a prerequisite, requiring substantial resources for high-quality segmentation masks and increasing the overall prediction time. To address these limitations, we propose CAP, a novel end-to-end one-stage framework that reimagines cell tracking by treating Cell as Point. Unlike traditional methods, CAP eliminates the need for explicit detection or segmentation, instead jointly tracking cells for sequences in one stage by leveraging the inherent correlations among their trajectories. This simplification reduces both labeling requirements and pipeline complexity. However, directly processing the entire sequence in one stage poses challenges related to data imbalance in capturing cell division events and long sequence inference. To solve these challenges, CAP introduces two key innovations: (1) adaptive event-guided (AEG) sampling, which prioritizes cell division events to mitigate the occurrence imbalance of cell events, and (2) the rolling-as-window (RAW) inference strategy, which ensures continuous and stable tracking of newly emerging cells over extended sequences. By removing the dependency on segmentation-based preprocessing while addressing the challenges of imbalanced occurrence of cell events and long-sequence tracking, CAP demonstrates promising cell tracking performance and is 8 to 32 times more efficient than existing methods. The code and model checkpoints will be available soon.
@@ 1 Introduction
Figure 1: Our proposed CAP is an end-to-end trained framework leveraging the idea of Cell as Point to facilitate tracking cells efficiently. As (a) illustrates, instead of previous work requiring segmentation (SEG) or detection (DET) as a prerequisite for final tracking (TRA), CAP is able to track all cells within sequence frames in one stage. (b) shows that CAP reduces the inference time ($2.9$s) by approximately $\mathbf{8}$ to $\mathbf{32}$ times compared to previous works while maintaining high tracking performance of $0.93$. (c) demonstrates the cell tracking result predicted by CAP.
@@ 2.1 Cell Tracking
Cell tracking is a crucial task in biological imaging that typically involves the identification and temporal association of cell instances across image frames. Traditionally, this process includes two main steps: cell segmentation (detection) [1, 85, 60, 83] and cell linking [18]. The segmentation step identifies individual cells in each frame, while the linking step establishes temporal associations between these cells across frames, sometimes capturing complex dynamics such as cell division and migration [54, 57]. Nowadays, instead of traditional approaches such as Nearest Neighbors linking [73], Bayesian filters [72, 34, 16, 33, 87, 35, 79], Viterbi [56, 55], and graph-based matching [5, 78, 3, 53], the advent of deep learning has transformed the landscape of cell tracking, leveraging advanced neural architectures for segmentation and tracking tasks. Deep learning approaches often employ complex models, e.g., CNN, RNN, GNN, and transformer, to combine multiple stages like detection [12, 10, 11, 50, 94, 30, 39, 61, 84, 31, 32, 40, 45], segmentation [88, 65, 9, 2, 54, 74, 7, 8, 70, 63, 17, 13, 62, 14, 77, 26, 64], and tracking into the frameworks. Several contemporary methods [68, 66, 30] aim to integrate contextual cell data into tracking models. ConvGRU [66] is introduced by Payer et al., capturing local cellular and inter-frame features. However, this method requires extensive annotated training data for all cell regions, making it labor-intensive. Motion and Position Map (MPM) is capable of representing detection and association at once, which is proposed by Hayashida et al. [31], marking an advancement in simplifying the pipeline for cell tracking. However, it still requires multiple stages, first generating the MPM followed by subsequent detection and tracking procedures.
@@ 2.2 Point Tracking
3 Methodology
3.1 Preliminaries
Recurrent All-Pairs Field Transforms (RAFT) [76] embedded in CoTracker [43] can be conceptualized as incrementally computing the similarity between all pairs of pixel points across frames in the sequence [29], enabling the model to handle large displacements and complex motion patterns [71]. In the field of cell tracking, this design promotes the cell point’s trajectories in a frame sequence constructed by iteratively updating the flow field using a recurrent unit that performs lookups on a 4D correlation volume.
4D Correlation Volume.
Given two feature maps $F_{1}\in\mathbb{R}^{H\times W\times dim}$ and $F_{2}\in\mathbb{R}^{H\times W\times dim}$ with output dimdim-dimensions, which are features extracted from two input images $I_{1}$ and $I_{2}$, by the feature extractor $g_{\phi}$: $F_{1}=g_{\phi}(I_{1})$ and $F_{2}=g_{\phi}(I_{2})$. RAFT computes a correlation volume $C$ that captures the similarity between all pairs of pixels across the two images. The correlation between a pixel at position $(i,j)$ in $F_{1}$ and a pixel at position $(k,l)$ in $F_{2}$ is calculated as the dot product of their feature vectors:
$C_{ijkl}$ $=\langle F_{1}(i,j),F_{2}(k,l)\rangle=\sum_{h}F_{1}(i,j,h)\cdot F_{2}(k,l,h),$ (1)
which is a 4D tensor of dimensions $H\times W\times H\times W$, where $H$ and $W$ are respectively the height and width of feature maps $F$s. To capture multi-scale feature information for large displacement, RAFT constructs a correlation pyramid by applying average pooling to the last two dimensions of the 4D correlation volume. Specifically, this pooling is performed with kernel sizes $k$ of 11 22, 44, and $8$, and a set of multi-scale correlation volumes $C^{k}=\{C^{1},C^{2},C^{3},C^{4}\}$ is generated. Thus, each volume $C^{k}$ in the pyramid has dimensions $H\times W\times\frac{H}{2^{k}}\times\frac{W}{2^{k}}$.
Iterative Updates.
The update operator estimates a sequence of flow estimations $\{f_{1},\dots,f_{n}\}$ with initializing the start point $f_{0}=0$. For each iteration, it produces $\Delta f$ as the direction that should be applied to the current estimation: $f_{k+1}=\Delta f+f_{k+1}$. Specifically, given the current flow estimation $f_{k}$, the correlation volume pyramid is retrieved and undergoes processing through two convolutional layers to extract relevant information. The updated information is comprised of correlation, flow, and context features and is updated over time as a feature map concatenation.
3.2 Proposed Method: Cell as Point (CAP)
Figure 2: Cell Point Trajectory and Visibility. M, D1, and D2 denote mother cell, daughter cell #1, and daughter cell #2. (x,y) in $L_{t}$ represent the location coordinates, and 0 or 11 in $V_{t}$ respectively represent non-existing and existing cells. Three valid trajectory and visibility status: (a) M has not divided; (b) M has divided, and only D1 occurs in the $t$-th frame; (c) M has divided, and both D1 and D2 occur in the $t$-th frame.
@@ 3.2 Proposed Method: Cell as Point (CAP)
3.2.1 Adaptive Event-Guided (AEG) Sampling
@@
3.2.2 Cell Joint Tracking
@@
To initialize the tracking feature $F$ for the part of input tokens $G$, we first generate image features (feature map), $\phi(I_{t})$, for each frame in the training sequence, from which the tracking features, $F$, are distilled according to cell point locations. Specifically, the feature extractor is a convolutional neural network (CNN) trained end-to-end in the framework for computational efficiency. Leveraging the CNN feature extractor, the feature maps with dimdim dimensions are generated from $I_{t}$, $\phi(I_{t})\in\mathbb{R}^{dim\times\left(\left\lfloor\frac{H}{k}\right\rfloor\times\left\lfloor\frac{W}{k}\right\rfloor\right)}$, where $k=4$ for computational efficiency, and the features are downscaled by average pooling with strides $s$ for multi-scaled features. The tracking features $F^{i}_{t=1}$ for $i$ are captured by bilinear interpolation from the location of cell points on the initial frame $L_{t=1}$ on the feature map $\phi(I_{t=1})$. The tracking feature vector $F^{i}_{t}\in\mathbb{R}^{dim}$ is initialized for the transformer input by broadcasting the tracking features in $t=1$ frame for all $T_{s}$ frames, which is refined and updated with $M$ iterations.
@@
[…]
@@
3.2.3 Rolling-as-Window (RAW) Inference
In cell tracking, the inference stage is crucial for producing the final tracking mask and establishing cell lineage relationships by analyzing predicted cell locations and division events. Existing works [74, 70, 84, 54, 8, 52] often customize inference and post-processing stages with different configurations for each dataset, adapting to variations in sequence length. This data-specific tailoring limits their generalizability and broader applicability to other datasets. In contrast, our proposed Rolling-As-Window (RAW) inference approach in CAP simplifies and unifies these processes while maintaining effectiveness across varying sequence lengths, enabling high-quality inference regardless of computational constraints. Specifically, a consistent and high-quality inference result is achieved through a rolling loop, which sequentially extracts yet-to-be-inferred clips from the full sequence, processing them within a defined window for inference.
@@ 4.1 Experimental Setup
4.1.1 Datasets
@@
We evaluate our proposed framework on DynamicNuclearNet Tracking in DeepCell [59] public database, HeLa, PC-3, 3T3, and RAW264 datasets divided by specimen, and ISBI cell tracking challenge (ISBI CTC) [57, 80] public datasets, PhC-C2DH-U373 (U373), Fluo-N2DH-GOWT1 (GOWT1), and Fluo-C2DL-Huh7 (Huh7). These temporal data used for cell tracking, shot by different devices in various durations on diverse specimens, exhibit diverse morphologies and sequence characteristics.
Figure 4: Visualization of the DeepCell Dataset. The datasets HeLa, PC-3, 3T3, and RAW264 have different cell densities, sizes, and luminances.
DeepCell.
DeepCell database consists of HeLa, PC-3, 3T3, and RAW264 datasets, which respectively contains cell types HeLa-S3, PC-3, NIH-3T3, and RAW 264.7, visualized in Fig. 4. These four cell types are mammalian cell lines that were acquired from the American Type Culture Collection [59]. The cells are cultured in Dulbecco’s modified Eagle’s medium (RAW 264.7 and NIH-3T3) or F-12K medium (Hela-S3 and PC-3) supplemented with the necessary drugs, e.g., L-glutamine, penicillin, for all other cells. DeepCell contains $96$ sequences in total: $45$ for HeLa, $7$ for PC-3, $10$ for 3T3, and $34$ for RAW264. Among them, there are $5$, 22, 11, and $3$ sequences for datasets HeLa, PC-3, 3T3, and RAW264 for evaluating the training outcomes of the remaining sequences, respectively. The detailed information on each cell type is illustrated below: